OpenStudy (anonymous):

how do you solve x^3 - 8 ??????

OpenStudy (shamil98):

x^3 - 8 = 0 x^3 = 8 take the cubed root of both sides x = 2

OpenStudy (anonymous):

Thank You! so when you are taking the cubed root out of 8 all you're doing is dividing it by 3?

OpenStudy (shamil98):

no. \[\huge \sqrt[3]{8} = 2\]

OpenStudy (shamil98):

the cubed root of a number means what value to the power of 3 makes this value?

OpenStudy (shamil98):

for example \[\huge \sqrt[3]{64} = 4 , 4^3 = 64\]

OpenStudy (anonymous):

okay..can you tell me how to type that into a graphing calculator if possible??

OpenStudy (anonymous):

or is it just something where i have to try my luck to find the answer?